OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 2i

OpenStudy (anonymous):

zeros are 5,-3,-1+2 i,-1-2 i (because imaginary roots always occur in pairs) required polynomial =(x-5)(x+3)(x+1-2i)(x+1+2i)=? [{x+1-2i}{x+1+2i}={(x+1)-2i}{(x+1)+2i}=(x+1)^2-(2i)^2=x^2+2x+1-4i^2 =x^2+2x+1+4=x^2+2x+5} required polynomial=(x-5)(x+3)(x^2+2x+5) simplify .